# What Linguistic Features and Languages are Important in LLM Translation?

### 21 Feb 2024 v1

@@ Abstract
Large Language Models (LLMs) demonstrate strong capability across multiple tasks, including machine translation. Our study focuses on evaluating Llama2’s machine translation capabilities and exploring how translation depends on languages in its training data. Our experiments show that the 7B Llama2 model yields above 10 BLEU score for all languages it has seen, but not always for languages it has not seen. Most gains for those unseen languages are observed the most with the model scale compared to using chat versions or adding shot count. Furthermore, our linguistic distance analysis reveals that syntactic similarity is not always the primary linguistic factor in determining translation quality. Interestingly, we discovered that under specific circumstances, some languages, despite having significantly less training data than English, exhibit strong correlations comparable to English. Our discoveries here give new perspectives for the current landscape of LLMs, raising the possibility that LLMs centered around languages other than English may offer a more effective foundation for a multilingual model.
@@
What Linguistic Features and Languages are Important in LLM Translation?
@@ 1. Introduction
Large Language Models (LLMs) and their utilizations have been a popular research topic in Natural Language Processing (NLP) due to their remarkable performance on various tasks Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b) and machine translation is no exception. Extensive evaluations Robinson et al. (2023); Hendy et al. (2023) on machine translation of the prominent GPT OpenAI (2023) models have shown that they exhibit a promising capability for high-resource machine translation although sometimes not for underrepresented languages Robinson et al. (2023); Hendy et al. (2023); Stap and Araabi (2023); Kadaoui et al. (2023).
@@
A popular approach for low-resource NLP is to leverage other higher-resourced languages. These approaches include using them as pivot Xia et al. (2019), in transfer learning Gu et al. (2018); Nguyen and Chiang (2017) and for joint training Neubig and Hu (2018); Johnson et al. (2017). Performance gains from such methods indicate a strong influence of their presence in the training data. Given that including related languages alongside the low-resource language can often improve performance Xia et al. (2019); Poncelas and Effendi (2022); Gu et al. (2018); Nguyen and Chiang (2017); Neubig and Hu (2018); Johnson et al. (2017), it makes sense to evaluate the degree of proximity between these languages, which can be done using the vectors from the URIEL typological database Littell et al. (2017). The utilization of the URIEL database has facilitated a more guided approach to exploring the incorporation of multiple languages into a single model by leveraging linguistically-aware feature vectors from which linguistic distances can be computed. These feature vectors have been utilized by previous works in various ways including determining which language to use as transfer or pivot language Lin et al. (2019); Nambi et al. (2023) and measuring language diversity Ruder et al. (2021).
It is ideal for an LLM to be as transparent as possible to conduct a comprehensive performance evaluation. This includes details such as the specific languages, the quantity of data for each language, and the nature of the data used for training. A significant constraint in studies that assess the GPT model series Hendy et al. (2023); Robinson et al. (2023) is the fact that these models are proprietary, closed-source systems. This presents a fundamental challenge as it remains unclear which languages are included in the models. On the other hand, determining the significance of language features poses a non-trivial challenge due to its many dimensions. For instance, the URIEL database vectors encompass several linguistic features, which include FEATURAL, GENETIC, GEOGRAPHICAL, PHONOLOGICAL, INVENTORY, and SYNTACTIC distances.
In this work we are evaluating Llama2 Touvron et al. (2023b) in machine translation to highlight its multilingual capability with an emphasis on languages the model has and has not seen during training. We analyze the results against linguistic feature distances to gather insight about which features and languages in the training data are important in LLM machine translation. In summary, our contributions are as follows:
@@
We provide comprehensive evaluation on 41 languages (15 of which not seen during training) for Llama2 model families.
@@
We reveal that increasing model parameters is potentially more effective in improving translation over instruction tuning and few-shot learning.
@@
We demonstrate that syntactic similarity is not always the most important linguistic feature nor are English features always the strongest correlating factors with machine translation scores.
@@ 2. Methodology
2.1. Language Similarity Metric Evaluation
| Language | ISO 639-3 | Writing System | Genus |
Afrikaans |
afr |
@@ 2.1. Language Similarity Metric Evaluation
sin |
Sinhala |
Indic |
Tagalog |
tgl |
@@
Celtic |
We experiment with languages reported in the training data of Llama2 Touvron et al. (2023b), the list of which and their respective ISO 639-3 codes can be found in Table 3. We refer to this set of languages as inllama. We also pick 15 languages not reported in the training data which we will refer to as outllama, presented in Table 1. It is important to highlight that languages not explicitly mentioned in Llama2 might still be present in the training data, albeit at a minuscule proportion of less than 0.005% Touvron et al. (2023b). Languages in outllama cover various language genera and writing systems. The machine translation evaluation is conducted using the FLORES-200 Guzmán et al. (2019) benchmark as it provides support for numerous low-resource languages, we exclude into-English translation directions to mitigate the risk of potential data leakage, given that the benchmark relies on Wikipedia as its English data source. We also exclude zero-shot translation as LLMs often get the language wrong in this prompting setup as reported by Robinson et al. (2023). We use a statistical machine translation metric, the BLEU Papineni et al. (2002) score and a model-based machine translation metric when applicable, the default COMET-22 Rei et al. (2022) model.
@@
The machine translation evaluation begins with oneshot translations for both languages in inllama and outllama using the 7B model. For every language covered in this experimental setup, distances will be calculated from every language in inllama on five dimensions, namely GENETIC, GEOGRAPHICAL, INVENTORY, PHONOLOGY, and SYNTACTIC as per the URIEL typological database Littell et al. (2017). We leave out the FEATURAL distances to give each of the dimensions more focus as FEATURAL distances are combinations of all the other feature distances111For more detailed explanation of these distances, consult https://www.cs.cmu.edu/~dmortens/projects/7_project/. We calculate the Pearson correlation for each linguistic distance feature against the translation scores. We identify languages in outllama that yields under 10 BLEU score and consider them truly unseen languages. For truly unseen languages, we experiment further with model scale, chat version, and adding the shot-count (random 5 shots). For translation with chat models, we use the prompt by Robinson et al. (2023) which follows the recommendation of Gao et al. (2023) to design prompts for translation using instruction-tuned models. The prompts used on our experiments are given in Table 2
| Language | ISO 639-3 | BLEU | COMET-22 |
German |
deu |
33.68 |
0.83 |
Swedish |
swe |
37.71 |
0.87 |
Dutch |
nld |
27.45 |
0.84 |
Norwegian |
nor |
29.54 |
0.86 |
Danish |
dan |
36.21 |
0.86 |
French |
fra |
42.4 |
0.84 |
Spanish |
spa |
28.54 |
0.84 |
Italian |
ita |
28.78 |
0.85 |
Portuguese |
por |
43.21 |
0.87 |
Catalan |
cat |
35.92 |
0.84 |
Romanian |
ron |
31.58 |
0.84 |
Russian |
rus |
28.21 |
0.85 |
Polish |
pol |
22.34 |
0.83 |
Ukrainian |
ukr |
26.03 |
0.83 |
Serbian |
srp |
23.96 |
0.81 |
Czech |
ces |
24.94 |
0.82 |
Bulgarian |
bul |
29.57 |
0.83 |
Croatian |
hrv |
21.3 |
0.81 |
Slovenian |
slv |
19.51 |
0.77 |
Chinese |
zho |
19.79 |
0.82 |
Japanese |
jpn |
17.02 |
0.84 |
Vietnamese |
vie |
28.77 |
0.82 |
Korean |
kor |
11.08 |
0.78 |
Indonesian |
ind |
31.15 |
0.86 |
Finnish |
fin |
18.08 |
0.82 |
Hungarian |
hun |
18.4 |
0.78 |
@@ 3. Results and Analysis
3.1. Llama2 Translation Results
One-shot 7B Llama2 translation results are presented in Table 3 and Table 4. From Table 3, we observe that none of the languages included in inllama produce a BLEU score below 10. This suggests that we can reasonably assume that Llama2 is capable of translating into all the languages it has encountered during training. However, many languages in outllama yields a BLEU score under 10, this is expected as Llama2 is presumably unfamiliar with these languages. We move forward with languages in outllama that yields a BLEU score below 10 and experiment with other variations of Llama2. We explore the effect of scale, chat version, and adding shot count and present the results on Table 5. Due to our limited compute resources we excluded the 70B and 70B-chat versions of Llama2.
Model scale enhances translation ability, but not always for instruction tuning and not so much for adding shot count.
Our experiment results presented in Table 5 demonstrate that the 13B versions of Llama2 outperforms the smaller 7B versions for all truly unseen languages. However, larger models do not seem to yield the same number of gains for every language. In best cases, 13B models increase on average as high as 2.53 BLEU with a standard deviation of 1.64. We also observed that instruction-tuning through chat versions improves translations for some languages over models with the same scale and shot count. This is observed best on Igbo and Javanese, with gains as much as 3.16 and 2.87 respectively, and worst on Tagalog, which performs worse on chat models with a decrease as big as 2.64. Adding the shot count generally improves performance although it is less drastic as model scale and instruction-tuned setups with a mean increase of 0.47 and 0.08 for non-chat and chat Llama-13B respectively. While previous three factors appears to greatly enhance Llama2’s capacity to translate into some languages, there are languages where the prospects are quite limited. For instance, for Sinhala and Tamil, scaling the model results in gains only less than 1 BLEU score and neither does using chat models or adding shot count.
@@ 3.1. Llama2 Translation Results
3.2. Language Proximity Analysis
The translation results using Llama2 7B with the one-shot prompting setup presented on Table 3 and Table 4 are used for linguistic proximity analysis. We first retrieve pre-computed distances222http://www.cs.cmu.edu/~aanastas/files/distances.zip from the URIEL database and retrieve only the distances between the languages we are translating into and the languages reported in Llama2. Self or identity distances e.g. Igbo distance to Igbo are excluded in the Pearson correlation calculation. We present our linguistic proximity analysis as heatmaps presented in Figure 1 and 2 for correlations with BLEU and COMET-22 respectively. The numbers presented are correlations between each translation language’s distance from each language in Llama2 for each feature in the URIEL database with the machine translation (MT) scores. We create several different heatmaps according to the translation languages considered. Heatmaps created are for inllama languages only, outllama languages only, both inllama and outllama languages, only Germanic languages, only Romance languages, only Slavic languages, and for languages belonging to other genus. It is important to highlight that a negative correlation implies that the closer a language is to a particular linguistic feature, the higher the MT score is likely to be and vice versa. We observe that MT score of a language correlate with Wikipedia article count of the language333Retrieved from https://meta.wikimedia.org/wiki/List_of_Wikipedias_by_language_group on October 2023, as high as 0.64 Pearson correlation with BLEU and 0.55 Pearson correlation with COMET-22.
Syntactic distance is not always the most important feature.
When including every inllama languages, BLEU and COMET-22 scores show consistently strong correlations with syntactic features, especially with Germanic and Romance languages. This finding may not be particularly groundbreaking, as we already understand that the languages in inllama predominantly belong to these language genus. However, when considering only outllama languages, translation quality seem to depend more on genetic and phonological distances although it is often inconsistent whether it is a positive or negative correlation. Syntactic distances on the other hand, show almost no correlation to MT scores except for certain languages like Norwegian and Catalan, which consistently display a strong tendency to correlate with both MT metrics. Correlation with features other than syntactic is also observed when considering languages in other genus, in which proximity with the inventory feature of Vietnamese, Dutch, German, and French are shown to correlate with COMET-22 scores.

### 05 Apr 2024 v2

# Could We Have Had Better Multilingual LLMs If English Was Not the Central Language?

@@ Abstract
Large Language Models (LLMs) demonstrate strong machine translation capabilities on languages they are trained on. However, the impact of factors beyond training data size on translation performance remains a topic of debate, especially concerning languages not directly encountered during training. Our study delves into Llama2’s translation capabilities. By modeling a linear relationship between linguistic feature distances and machine translation scores, we ask ourselves if there are potentially better central languages for LLMs other than English. Our experiments show that the 7B Llama2 model yields above 10 BLEU when translating into all languages it has seen, which rarely happens for languages it has not seen. Most translation improvements into unseen languages come from scaling up the model size rather than instruction tuning or increasing shot count. Furthermore, our correlation analysis reveals that syntactic similarity is not the only linguistic factor that strongly correlates with machine translation scores. Interestingly, we discovered that under specific circumstances, some languages (e.g. Swedish, Catalan), despite having significantly less training data, exhibit comparable correlation levels to English. These insights challenge the prevailing landscape of LLMs, suggesting that models centered around languages other than English could provide a more efficient foundation for multilingual applications.
@@
Could We Have Had Better Multilingual LLMs If English Was Not the Central Language?
@@ 1. Introduction
Large Language Models (LLMs) have been a popular research topic in Natural Language Processing (NLP) due to their remarkable performance on various tasks including machine translation Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b). Extensive evaluations on machine translation of the popular GPT model family OpenAI (2023) have suggested that they can translate high-resource languages Robinson et al. (2023); Hendy et al. (2023). However, it is rarely the case for low-resource or underrepresented languages Robinson et al. (2023); Hendy et al. (2023); Stap and Araabi (2023); Kadaoui et al. (2023).
A straightforward approach for the lack of training data in low-resource translation is to collect more labeled data. However, investing in data creation is nontrivial as it comes with challenges, including the cost of such endeavors. For example, Aji et al. (2022) described the absence of Wikipedia articles on Indonesian regional languages and the challenges of labeled data collection for them, which includes the lack of speakers, the diversity of dialects, and the lack of a writing standard. In addition, training large language models on more data brings environmental consequences Strubell et al. (2019). In the long run, more training data may require longer GPU compute hours, which will release more greenhouse gas emissions.
@@
Aside from data creation, other techniques are often employed as an alternative. A popular approach for multilingual or low-resource NLP is to leverage other languages to benefit from cross-lingual transfer. These approaches include using them as pivot Wijaya et al. (2017); Xia et al. (2019), transfer learning Gu et al. (2018); Nguyen and Chiang (2017), and joint training Neubig and Hu (2018); Johnson et al. (2017). Improvements from such methods indicate a strong influence of the presence of other languages in the training data. Given that including related languages alongside the low-resource language can improve performance Xia et al. (2019); Poncelas and Effendi (2022); Gu et al. (2018); Nguyen and Chiang (2017); Neubig and Hu (2018); Johnson et al. (2017), it is beneficial to include proximity measurements between these languages on evaluations, which can be done using the vectors from the URIEL database Littell et al. (2017). The utilization of the URIEL database has made evaluating multiple languages more explainable by leveraging linguistically aware feature vectors from which linguistic distances can be computed. These vectors have been utilized by previous works in various ways including determining which language to use as transfer or pivot language Lin et al. (2019); Nambi et al. (2023) and measuring language diversity Ruder et al. (2021).
It has been established that there are benefits to using other languages in the training process. However, multilingual labeled data creation is challenging. In this paper, we aim to provide hints to narrow down future data collection strategies by evaluating an existing LLM family. A constraint in previous studies that assess the GPT model series Hendy et al. (2023); Robinson et al. (2023) has been the fact that these models are proprietary, closed-source systems that do not disclose information regarding their training data. This presents a challenge as it remains unclear which languages are included in the training of the models. On the other hand, open-source LLMs such as Meta’s Llama2 Touvron et al. (2023b), is more transparent about its training process, including the languages that are included in its training data. This makes the model more suitable as a subject for our evaluation.
In this work, we are evaluating Llama2 Touvron et al. (2023b) for machine translation to highlight its multilingual capability in languages it has or has not seen during training. We also model a linear relationship (through correlation scores) between the linguistic feature distances and the translation metrics and use these scores as a basis for language importance analysis. The goal of the analysis is to narrow down the data investment effort by shedding light on which language(s) may improve the translation of other languages when included in the training data. An efficient data collection strategy will result in future multilingual LLMs that can be trained and deployed more efficiently, thus promoting sustainability. In summary, our contributions are as follows:
@@
We evaluate Llama2 and provide machine translation scores of this model for 41 languages, 15 of which were not seen during its training.
@@
We reveal that increasing model parameters is more effective in improving translation over instruction tuning and few-shot learning.
@@
Our research reveals that syntactic similarity between languages is not the only linguistic aspect that is strongly linked to machine translation performance. Surprisingly, these strong correlations between linguistic feature distances and machine translation performances extend beyond English and hold true across various languages, therefore opening up the possibility of other better central languages for multilingual LMs
@@ 2. Methodology
2.1. Machine Translation Evaluation
We experiment with languages reported in the training data of Llama2 Touvron et al. (2023b), the list of which and their respective ISO 639-3 codes can be found in Table 1. We refer to this set of languages as inllama. We also pick 15 languages not reported in the training data which we will refer to as outllama, presented in Table 2. It is important to highlight that languages not explicitly mentioned in Llama2 might still be present in the training data, albeit at a minuscule proportion of less than 0.005% of its training data Touvron et al. (2023b). Languages in outllama cover various language genera and writing systems. The machine translation evaluation is conducted using the FLORES-200 Guzmán et al. (2019) benchmark as it is available for numerous low-resource languages. We exclude XEnglish translation directions to mitigate the risk of potential data leakage, given that FLORES-200 uses Wikipedia for its English sentences. We also exclude zero-shot translation as LLMs often get the language wrong in this prompting setup as reported by Robinson et al. (2023). We measure translation quality using machine translation scores. Translation quality is measured with the BLEU score Papineni et al. (2002) and a model-based machine translation metric (COMET-22 Rei et al. (2022)) where applicable. COMET-22 is used to compensate for the drawbacks of BLEU and vice-versa.
| Language | Genus | BLEU | COMET-22 |
| German (deu) |
Germanic |
33.68 |
0.83 |
| Swedish (swe) |
Germanic |
37.71 |
0.87 |
| Dutch (nld) |
Germanic |
27.45 |
0.84 |
| Norwegian (nor) |
Germanic |
29.54 |
0.86 |
| Danish (dan) |
Germanic |
36.21 |
0.86 |
| French (fra) |
Romance |
42.4 |
0.84 |
| Spanish (spa) |
Romance |
28.54 |
0.84 |
| Italian (ita) |
Romance |
28.78 |
0.85 |
| Portuguese (por) |
Romance |
43.21 |
0.87 |
| Catalan (cat) |
Romance |
35.92 |
0.84 |
| Romanian (ron) |
Romance |
31.58 |
0.84 |
| Russian (rus) |
Slavic |
28.21 |
0.85 |
| Polish (pol) |
Slavic |
22.34 |
0.83 |
| Ukrainian (ukr) |
Slavic |
26.03 |
0.83 |
| Serbian (srp) |
Slavic |
23.96 |
0.81 |
| Czech (ces) |
Slavic |
24.94 |
0.82 |
| Bulgarian (bul) |
Slavic |
29.57 |
0.83 |
| Croatian (hrv) |
Slavic |
21.3 |
0.81 |
| Slovenian (slv) |
Slavic |
19.51 |
0.77 |
| Chinese (zho) |
Chinese |
19.79 |
0.82 |
| Japanese (jpn) |
Japanese |
17.02 |
0.84 |
| Vietnamese (vie) |
Vietic |
28.77 |
0.82 |
| Korean (kor) |
Korean |
11.08 |
0.78 |
| Indonesian (ind) |
Malayo-Sumbawan |
31.15 |
0.86 |
| Finnish (fin) |
Finnic |
18.08 |
0.82 |
| Hungarian (hun) |
Ugric |
18.4 |
0.78 |
We aim to experiment with open-source LLMs that replicate proprietary models such as ChatGPT OpenAI (2023) in terms of usability and safety. At the time the Llama2 model was released and the experiment design for this paper was constructed, none of the open-source models are suitable substitutes for production models as they may not have been aligned to match human preferences and there may be a performance gap Touvron et al. (2023b). On account of this, we decided to move forward only with the Llama2 model family. The machine translation evaluation begins with one-shot translations for both languages in inllama and outllama using the vanilla 7B model. From this experiment, we categorize languages that yield under 10 BLEU as unlearned languages111Based on ”Almost useless” interpretation from https://cloud.google.com/translate/automl/docs/evaluate. For the unlearned languages, we experiment further with model scale, chat version, and adding the shot count to maximize the potential of in-context learning. Our choice of randomly picking 5 shots from the validation set of FLORES-200 is motivated by the experimental setup used by Hendy et al. (2023) which states that increasing beyond 5 shots does not result in meaningful improvement and shows that selected quality shots do not always improve more than 1 BLEU compared to random selections for GPT (text-davinci-003) model. For translation with chat models, we use the prompt by Robinson et al. (2023) which follows the recommendation of Gao et al. (2023) for designing prompts for translation using instruction-tuned models. The prompts used in our experiments are given in Table 3
| Language | Genus | Writing System |
| Afrikaans (afr) |
Germanic |
@@ 2.1. Machine Translation Evaluation
| Tagalog (tgl) |
Greater Central Philippine |
@@
2.2. Correlation Score Analysis
We consider several language subsets. For every language subset, we calculate the Pearson correlation score between the linguistic similarity scores of each language in the subset to a language in inllama and their respective translation scores. We assume that a certain language is important if we observe a positive correlation. For example, consider the language A and the language subset {B, C, D, E}. When the similarity of A with each language in the subset {B, C, D, E} and the respective machine translation scores for {B, C, D, E} exhibit a positive correlation, i.e. the closer they are to A the better their machine translation scores, A is deemed as a valuable language and is therefore hypothesized to be more optimal for a central language when developing multilingual language models. A is checked for each language in inllama. Similarity scores are calculated on five dimensions: GENETIC, GEOGRAPHICAL, INVENTORY, PHONOLOGY, and SYNTACTIC as per the URIEL typological database Littell et al. (2017). We exclude the FEATURAL distances to focus on each dimension as FEATURAL distances are combinations of all the other feature distances222For more detailed explanation of these distances, consult https://www.cs.cmu.edu/~dmortens/projects/7_project/. Language subsets considered are inllama languages only, outllama languages only, both inllama and outllama languages, only Germanic languages, only Romance languages, only Slavic languages, and languages belonging to Other genera.
@@ 3. Results and Analysis
3.1. Machine Translation Evaluation Results
One-shot 7B Llama2 translation results are presented in Table 1 and Table 4. From Table 1, we observe that none of the languages included in inllama produce a BLEU score below 10. This suggests that we can reasonably assume that Llama2 is capable of translating into all the languages it has encountered during training. However, many languages in outllama yield a BLEU score under 10, this is expected as Llama2 is presumably unfamiliar with these languages. On the other hand, we hypothesize that there are two possibilities for the high-performing outllama languages; (1) those languages are indeed included in the training data i.e. included in the 0.005% of the training data, or (2) similar languages in inllama indeed boosted their performance.
We move forward with languages in outllama that yield a BLEU score below 10 and experiment with other variations of Llama2. We explore the effect of scale, chat version, and adding shot count and present the results in Table 5. Due to our limited compute resources we excluded the 70B and 70B-chat versions of Llama2.
Scaling up the model enhances translation ability. However, improvements from instruction-tuning and adding shot count remain inconclusive.
Results presented in Table 5 demonstrate that the 13B versions of Llama2 outperform the smaller 7B versions for all unlearned languages. However, larger models do not seem to yield the same number of gains for every language. In best cases, 13B models increase on average as high as 2.53 BLEU with a standard deviation of 1.64. For instruction-tuning (chat) models, we observed both performance increase and decrease. The best improvements are observed in Igbo and Javanese, which improves as much as 3.16 and 2.87 respectively, and a decrease is observed in Tagalog, which performs worse on chat models with a decrease as severe as 2.64. Adding the shot count generally improves performance although it is less drastic than model scale and instruction-tuning with a mean increase of 0.47 and 0.08 for non-chat and chat Llama-13B respectively. While these model variations appear to enhance Llama2’s capacity to translate into some languages greatly, there are languages where the prospects are limited. For instance, for Sinhala and Tamil, scaling up the model/adding shot count/using chat models results in less than 1 BLEU score increase.
@@ 3.1. Machine Translation Evaluation Results
3.2. Language Importance Analysis
We use the results from Table 1 and Table 4 for the linguistic proximity analysis. We first retrieve pre-computed distances333http://www.cs.cmu.edu/~aanastas/files/distances.zip from the URIEL database and retrieve only the distances between the languages we are translating into and the languages reported in Llama2. Self or identity distances e.g. Igbo-to-Igbo distance are excluded in the Pearson correlation calculation. This correlation analysis aims to model the linear relationship between language proximity and machine translation scores to identify languages whose data may be beneficial for multilingual training.
We present our analysis as heatmaps in Figure 2 and 3 for correlations with BLEU and COMET-22 respectively. To help understand where each number came from in the heatmap, a scatter plot visualization for SYNTACTIC distance to Swedish for the combined inllama and outllama language subset against BLEU scores is presented in Figure 1 as an example. We create several different heatmaps according to the subset considered. It is important to highlight that distance is used as a similarity score. Therefore, a negative correlation between linguistic distance and MT scores would imply that the closer (i.e., the smaller the linguistic distance) a language is to this language, the higher the MT score is likely to be. In addition, since Wikipedia is a permanent fixture of LLMs’ training data, we observe that there is a positive correlation between MT scores and Wikipedia article counts444Retrieved from https://meta.wikimedia.org/wiki/List_of_Wikipedias_by_language_group on October 2023, as high as 0.64 using BLEU and 0.55 using COMET-22.
Syntactic similarity may be an important feature, but other linguistic dimensions can be too.
When including every language, i.e. the inllama and outllama subset, BLEU complemented with COMET-22 scores show consistently strong correlations with syntactic features, especially with Germanic and Romance languages. This finding may not be particularly groundbreaking, as we already understand that the languages in inllama predominantly belong to these language genus. However, when considering only outllama language subset, translation performance seems to have higher correlations (either positive or negative) with GENETIC and PHONOLOGICAL distances. When considering languages in outllama, only SYNTACTIC similarities to certain languages e.g. Norwegian and Catalan display a strong correlation with MT performances. Correlation with features other than SYNTACTIC is also observed when considering languages in other genera, in which the proximity with the INVENTORY feature of Vietnamese, Dutch, German, and French are shown to correlate with COMET-22 scores.